• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name gilman

    Log entry time 08:37:26 on August28,2006

    Entry number 183410

    keyword=owl shift summary 28 Aug 2006


    Started shift in kin #13: theta_left = 46 deg, theta_right = 24 deg,
    p0_left = 0.3604 GeV/c
    runs target raster I uA events
    3485/21760 4cm dummy on 7 1.5M/2.2M
    3486/21761 4cm dummy on 7 0.3M/0.5M MPS BCM problem starts
    3487/21762 4cm dummy on - junk, during MPS BCM problem
    3488/21763 4cm dummy on 7 junk, during beam recovery
    3489/21764 4cm dummy on 7 junk, capture section problems
    beam comes back and we have DAQ problems -- some junk runs ---
    3492/21767 4cm dummy on 7 0.5M/1.1M
    3493/21768 4cm dummy off 5 1M/1.9M
    3494/21769 4cm dummy off 5 1M/1.9M
    -- set p0 =0.3418 GeV/c --
    3495/21770 4 cm dummy off 5 0.5M/0.5M stop to reset PS-lefts
    3496/21771 4 cm dummy off 5 0.3M/0.3M stop to reset PS-lefts
    3497/21772 4 cm dummy off 5 1.0M/1,7M
    3498/21773 LD2 off 5 1.0M/1.1M
    3499/21774 LD2 off 5 1.0M/1,1M
    3500/21775 LD2 on 5 1.0M/1.1M
    3501/21776 LD2 on 5 1.0M/1.1M
    3502/21777 4 cm dummy on 5 1.0M/2.8M
    3503/21778 4 cm dummy on 5 1.0M/2.8M